OpenStudy (anonymous):

2x^2-5x-4=0 Using the quadratic formula

jimthompson5910 (jim_thompson5910):

\[\Large x = \frac{-b\pm\sqrt{b^2-4ac}}{2a}\] \[\Large x = \frac{-(-5)\pm\sqrt{(-5)^2-4(2)(-4)}}{2(2)}\] \[\Large x = \frac{5\pm\sqrt{25-(-32)}}{4}\] I'll let you finish